OpenStudy (abbycross167):

Can someone please check my question/answer? Please? Which values of x satisfy the inequality –(7/3)[x+(9/28)]<20? a. x < -8 25/28 b. x < -8 ¼ c. x > -8 25/28 d. x > -8 ¼

OpenStudy (abbycross167):

I think its d

OpenStudy (ineedhelplz):

Is that negative?

OpenStudy (abbycross167):

sir/ma'am?

OpenStudy (ineedhelplz):

–(7/3)[x+(9/28)]<20?

OpenStudy (ineedhelplz):

The beginning

OpenStudy (ineedhelplz):

Because if it is negative, I got C

jhonyy9 (jhonyy9):

@Ineedhelplz please step by step

jhonyy9 (jhonyy9):

7 9 - (---)( x + (---) ) < 20 3 28 @abbycross167 this is your exercise yes ?

OpenStudy (abbycross167):

yes sir

jhonyy9 (jhonyy9):

so like a first step i think will be right than you divide both sides by (-7/3) do you can make it ?

jhonyy9 (jhonyy9):

i like teach you step by step - ok. ?

OpenStudy (abbycross167):

yes sir/ma'am

jhonyy9 (jhonyy9):

so than from the left side will be elimineted the fraction (-7/3 ) and on the right side what will get ?

OpenStudy (abbycross167):

I'm not sure,

OpenStudy (abbycross167):

like do I eliminate 20? is that what you meant?

jhonyy9 (jhonyy9):

no - from the left side will be eliminated the -7/3 fraction on the right side you need dividing the 20 by (-7/3) how you make it ? 20 ------ = ? -7 ( --- ) 3

jhonyy9 (jhonyy9):

for more easy way you can rewrite it (20/1) ------ = ? (-7/3) hope you know how divide two fractions - yes ?

OpenStudy (abbycross167):

-8 4/7?

jhonyy9 (jhonyy9):

multiplie first by invers of second yes ?

jhonyy9 (jhonyy9):

try it please and see what will get

OpenStudy (abbycross167):

-7/3*20/1= -46 2/3 is that what you wanted me to do?

jhonyy9 (jhonyy9):

yes but like result will be more easy to continue your calculi than will get fractionaly form without integer part - ok. ?

jhonyy9 (jhonyy9):

so -140/3 ok.?

jhonyy9 (jhonyy9):

but this not is right i have said multiplie first fraction with second inversed

jhonyy9 (jhonyy9):

so will be 20/1 *(-3/7) = ?

jhonyy9 (jhonyy9):

so what will get ?

OpenStudy (abbycross167):

-8 4/7

jhonyy9 (jhonyy9):

-60/7 in the right side (x+(9/28))<(-60/7)

jhonyy9 (jhonyy9):

x+9/28 = ? how you add them ?

jhonyy9 (jhonyy9):

what will be the common denominator ?

jhonyy9 (jhonyy9):

denominator common wil be 28 than multiplie the x by 28 and will get (28x+9)/28 < (-60/7)

jhonyy9 (jhonyy9):

there are two fractions too so you need getting the denominator common for 28 and 7 what number will be the denominator common ?

jhonyy9 (jhonyy9):

are you here please ?

OpenStudy (abbycross167):

I'm sorry my Wifi messed up

jhonyy9 (jhonyy9):

for 28 and 7 the denominator common will be 28 because 7*4=28 so 28x+9 - 60 ------ < ----- 28 7 can you continue ?

jhonyy9 (jhonyy9):

with what number you need multiplie the right side than you know that the denominator common is 28 ?

jhonyy9 (jhonyy9):

by 28/7 = 4 than you multiplie the right side by 4 what wil get in numerator ?

jhonyy9 (jhonyy9):

4*(-60) = ?

OpenStudy (abbycross167):

-240

jhonyy9 (jhonyy9):

yes right so now you can rewrite them witout denominator - so how will be ?

jhonyy9 (jhonyy9):

on the left side 28x+9 < ?

OpenStudy (abbycross167):

7?

jhonyy9 (jhonyy9):

why 7 when you have wrote above -240 for the right side ?

OpenStudy (abbycross167):

oh sorry i forgot...

jhonyy9 (jhonyy9):

than will get 28x+9 < - 240 can you solve it now for x ?

jhonyy9 (jhonyy9):

subtract from both sides 9 and will get ?

OpenStudy (abbycross167):

-230?

jhonyy9 (jhonyy9):

28x +9 -9 < - 240 -9 do you can continue it now ?

OpenStudy (abbycross167):

28+<-249

jhonyy9 (jhonyy9):

28x<-249 what will be the value of x ?

OpenStudy (abbycross167):

-8 25/28 ?

jhonyy9 (jhonyy9):

so this is just what will be on the right side but on the left side what will be ?

jhonyy9 (jhonyy9):

x< yes ?

jhonyy9 (jhonyy9):

so than what will be the right choices from these ?

jhonyy9 (jhonyy9):

choice - sorry

jhonyy9 (jhonyy9):

d. b. a. c. what will be right ?

OpenStudy (abbycross167):

a?

jhonyy9 (jhonyy9):

are you sure ? you have got this result sure ?

jhonyy9 (jhonyy9):

yes this is sure the right answer hope you understand it all so was my pleasure good luck bye

OpenStudy (abbycross167):

thank you very much for helping me, you're a great help!